Find three consecutive integers such that the sum of the least integer and the middle integer is 46 more than the greatest integer.

Answer by checkley79(3341)

LET X, X+1 & X+2 BE THE 3 INTEGERS
X+(X+1)=(X+2)+46
2X+1=X+2+46
2X-X=2+46-1
X=47 ANS.
PROOF:
47+48=49+46
95=95